Please help me solve this problem.





The probability of an event occurring is 7/9. Find the odds against the event occurring


Answer by stanbon(57307)

The probability of an event occurring is 7/9. Find the odds against the event occurring
---
P(event NOT occurring) = 2/9
---
Odds against = P(not)/P(for) = (2/9)/(7/9) = 2:7
====================
Cheers,
Stan H.
==========